# Characterisation of birefringence inhomogeneity of KAGRA sapphire mirrors from transmitted wavefront error measurements

### v2

@@ Abstract
Cooling down test masses to cryogenic temperatures is a way to reduce the thermal noise of gravitational wave detectors. Crystalline materials are considered the most promising materials for fabricating cryogenic test masses and their coatings because of their excellent thermal and optical properties at low temperatures. However, birefringence owing to local impurities and inhomogeneities in the crystal can degrade the performance of the detector. The birefringence measurement or mapping over a two-dimensional area is thus important. This study describes a method for fast birefringence measurements of a large sample by simply combining a series of transmission wavefront error measurements using linearly polarised light with Fizeau interferometers. Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses with a diameter of 22 cm was fully reconstructed. The birefringence information was then used to calculate the transverse beam shapes of the light fields in orthogonal polarisation directions when passing through the substrate. The calculated beam shapes showed a good agreement with the in-situ measurements using the KAGRA interferometer. This technique is crucial for birefringence characterisation of test masses in future detectors, where even larger sizes are used.
@@ I Introduction
Thermal noise is one of the main noise sources for gravitational wave detectors [1, 2]. To reduce thermal noise, future detectors such as the LIGO Voyager [3] and Einstein Telescope [4, 5] are designed to operate at cryogenic temperatures. Cosmic Explorer [6] considers cryogenic operations for upgrades. KAGRA [7, 8, 9, 10] is the only detector operating at cryogenic temperatures and pioneers this technique among the current large-scale ground-based detectors. All these detectors use large, high-quality mirrors as their test masses, forming optical resonators to sense the tiny length fluctuations introduced by gravitational waves. Fused silica is commonly used to make test masses at room temperature; however, it is not compatible with low temperatures. Crystalline materials are considered to be the most promising materials for fabricating cryogenic test masses. Sapphire was selected as the material for KAGRA’s test masses [11] because of its excellent thermal and optical properties at low temperatures [12, 13]. However, recent studies have revealed that the birefringence of crystalline substrates and coatings can degrade detector performance [14, 15, 16, 17, 18, 19]. In this case, the birefringence fluctuates in magnitude and direction throughout the substrate of the mirror. Therefore, it is necessary to measure the position dependence of the birefringence over a large area to assess its impact on optical performance.
@@
There are several studies on two-dimensional birefringence measurements of transparent materials [14, 15, 20, 21, 22, 23, 24]. These measurements use either a single-beam scanning technique or expand the beam to a large aperture for imaging using a sensor array. The scanning method is rather time-consuming for large sample sizes, for example, mirrors with diameters usually greater than 20 cm for ground-based gravitational wave detectors, if one wants to obtain good spatial resolution [25, 26]. A phase-shifting method in digital interferometry has been proposed for fast 2D birefringence measurements with high accuracy [27, 28, 29]. Fizeau interferometry is among the most accurate methods for extracting the phase information of an optical component (reflection or transmission phase) over a large area. The Transmission Wavefront Error (TWE) is often used to measure the refractive index homogeneity of a substrate under test. However, in the presence of birefringence a simple TWE measurement does not reveal all the information about the substrate.
@@
In this study, we describe a method for combining a series of TWE map measurements using linearly polarised light in an appropriate way to fully reconstruct refractive index information, including birefringence. Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses (ITM) is extracted. This birefringence information was then used to calculate the transverse beam shape of the light fields in orthogonal polarisation directions when passing through the substrate. The calculated beam shapes were in good agreement with the in-situ beam shape measurements. The method of extracting birefringence information from TWE measurements with Fizeau interferometers is introduced in Section II. The birefringence characterisation of KAGRA’s current ITMs is described in Section III, and the conclusions are summarised in Section IV.
@@ II Generating a birefringence map with a Fizeau interferometer
Figure 1 shows a schematic diagram of the Fizeau interferometer. We analyse the interference pattern between the light reflected by the reference plane 1 ($E_{\mathrm{1}}$) and the light reflected by the reference plane 2 ($E_{\mathrm{2}}$) on the CCD. By comparing the results obtained with and without the sample inserted into the cavity formed by reference planes 1 and 2, we obtained information on the optical thickness of the sample. Because the optical thickness is the physical thickness multiplied by the index of refraction, by measuring the surface figure maps of S1 (surface with anti-reflection coating) and S2 (surface with highly reflective coating) precisely and subtracting the variation in the physical thickness of the sample, we can extract the fluctuation of the index of refraction in the substrate.
@@ II.1 Transmission of a linearly polarised light through a birefringent medium
where $M$ denotes the round-trip Jones matrix of the substrate. The indices of refraction for the extraordinary and ordinary axes are denoted as $n_{\mathrm{e}}$ and $n_{\mathrm{o}}$, respectively. Then, we can compute the round- trip phase changes of the optical fields polarised along the axes:
$\alpha_{\mathrm{e}}\equiv 2\pi\frac{2dn_{\mathrm{e}}}{\lambda},$ (3)
@@
Using $\alpha_{\mathrm{e}}$ and $\alpha_{\mathrm{o}}$, we define the common and differential phase changes as follows:
$\alpha_{+}\equiv\frac{\alpha_{\mathrm{e}}+\alpha_{\mathrm{o}}}{2},$ (5)
$\alpha_{-}\equiv\frac{\alpha_{\mathrm{e}}-\alpha_{\mathrm{o}}}{2}.$ (6)
@@ II.3 Extraction of $\alpha_{-}$ and $\theta$ from several TWE maps
We can see $\psi(\eta)$ as a function of $\eta$ with parameters $\theta$ and $\alpha_{-}$ determining its shape. Therefore, if we have several measurements of $\psi(\eta)$ with different values of $\eta$, we can use a non-linear fitting algorithm to fit the measured values with Equation (12) to obtain the best fit values of $\theta$ and $\alpha_{-}$.
@@
III Birefringence characterisation of KAGRA sapphire ITMs
@@
Sapphire is a uniaxial crystal with an optical axis called the c-axis. KAGRA’s sapphire test masses are fabricated such that the input surface is perpendicular to the c-axis. Light propagating in the test mass substrate travels along the c-axis, and theoretically, should experience no birefringence. However, local impurities and inhomogeneities in the crystal, as well as mechanical and thermal stress (e.g. due to the suspension system supporting the mirror), can cause depolarisation. These local defects distorted the polarisation distribution of the light field in the transverse plane. From Equation (2), we can see that the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ can be eliminated when the differential phase $\alpha_{-}$ equals 00. In this case, there was no depolarisation for the input polarisation. Gravitational wave detectors have very strict requirements on the TWE of light going through input test mass substrates. To achieve this requirement and minimise the influence of crystal inhomogeneity, when the birefringence was ignored, an ion beam figuring technique was used during the polishing process, in which the thickness of the substrate was precisely adjusted over the surface to ensure that the TWE rms was within a few nanometres [18]. However, in a birefringent medium, $\alpha_{+}$ and $\alpha_{-}$ in Equation (2) are both functions of thickness. This results in the consequence that $\alpha_{+}$ and $\alpha_{-}$ of the wavefront distortions due to inhomogeneity and birefringence cannot be minimised simultaneously by ajusting the thickness of the substrate. Wavefront errors and depolarisation are trade-offs in a birefringent medium.
@@ III.1 TWE of KAGRA sapphire test mass
The sapphire ITMs have already been installed in KAGRA, and direct birefringence mapping is impossible now. However, characterisation of the sapphire test masses was studied in [31]. The TWE maps of the ITMs were measured several times with linearly polarised light by rotating the sample at different angles, which was equivalent to the measurements obtained by rotating the input polarisation. The TWE maps were measured at the angles of 0, 45, 90 and 135. The methods described in Section II require a rigid angular control. However, the main motivation of the measurements in [31] was to see the stress induced change in $\alpha_{+}$, not $\alpha_{-}$, by different mounting orientations. Therefore, the rotation angle was not precisely controlled. To mitigate the problem, we compared the locations of the markers on the edge of the mirrors in different maps. In this way, we were able to recover the real rotation angle (Figure 3). The TWEs of the KAGRA current ITMs measured with linearly polarised light are around 5 to 6 times larger than the specifications because the vendor used circularly polarised light to measure the path length difference and calculate the removal from surface S2 during the ion beam figuring process (rms values are listed in Table 1). As shown in Figure 3, the 0 map and the 88 map have almost opposite values (the same is true for 49 and 130 maps). By summing the two maps, we can derive the TWE for circularly polarised light, which has small residuals.
@@ III.2 Constructing birefringence maps from TWE
The biggest challenge for the birefringence characterisation of KAGRA’s current ITMs is that the previous TWE measurement setup was not optimised for birefringence studies. According to Equation (20) in Section II.3, to extract the birefringence information, we need to subtract two pairs of TWE maps and then divide. In typical Fizeau interferometer measurements, low-order effects such as the piston and surface tilt are removed from the results. When doing measurements with different $\theta$, the distance and alignment between the sample and reference mirror have changed after rotating the sample. This implies that the difference between the two maps contains an unknown piston and tilt. The numerator and denominator in Equation (20), which contain these unknown terms, provide incorrect birefringence information.
The centre of the map also changed after the sample was rotated. However, we observed that the decentring was very small (less than 2 mm), and its effect was negligible compared with the effects of the piston and tilt. In addition, astigmatism was observed on the original TWE maps. The left plot in Figure 4 shows the TWE map after removing the curvature. A clear astigmatism term can be observed on the map plotted on the right. We observed that the rotation of this astigmatism did not change when the samples were rotated. Therefore, we believe that the astigmatism originates from the gravity-induced deformation of the sample and the astigmatism of the reference mirror used for the Fizeau interferometer. Astigmatism was individually removed from each TWE map.
@@ III.3 Comparison with in-situ measurements
The derived $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY are shown in Figure 10. From these maps, we can see that the birefringence in KAGRA’s current ITMs is inhomogeneous. If we compare these maps with the beam shape in p-polarisation shown in Figure 8, it can be found that the beam shape is mainly determined by the shape of $\theta$ map. This is because $\alpha_{-}$ is small in our case. The $\sin 2\theta$ term plays a more important part than $\sin\alpha_{-}$ in the amplitude of the field in p-polarisation from Equation (7).
@@ Acknowledgements.
We would like to thank Shinji Miyoki and Tomotada Akutsu at KAGRA for their support for the measurements. We also thank Masaki Ando, Kentaro Somiya, Kentaro Komori and Satoru Takano for insightful discussions. This work was supported by JSPS KAKENHI Grant No. JP20H05854, No. 22H01228, No. 23H01205, and by JST PRESTO Grant No. JPMJPR200B. KAGRA is supported by MEXT, JSPS Leading-edge Research Infrastructure Program, JSPS Grant-in-Aid for Specially Promoted Research 26000005, JSPS Grant-inAid for Scientific Research on Innovative Areas 2905: JP17H06358, JP17H06361 and JP17H06364, JSPS Core-to-Core Program A. Advanced Research Networks, JSPS Grantin-Aid for Scientific Research (S) 17H06133 and 20H05639 , JSPS Grant-in-Aid for Transformative Research Areas (A) 20A203: JP20H05854, the joint research program of the Institute for Cosmic Ray Research, University of Tokyo, National Research Foundation (NRF), Computing Infrastructure Project of Global Science experimental Data hub Center (GSDC) at KISTI, Korea Astronomy and Space Science Institute (KASI), and Ministry of Science and ICT (MSIT) in Korea, Academia Sinica (AS), AS Grid Center (ASGC) and the National Science and Technology Council (NSTC) in Taiwan under grants including the Rising Star Program and Science Vanguard Research Program, Advanced Technology Center (ATC) of NAOJ, and Mechanical Engineering Center of KEK.

### v3

@@ Abstract
Cooling down test masses to cryogenic temperatures is a way to reduce the thermal noise of gravitational wave detectors. Crystalline materials are considered the most promising materials for fabricating cryogenic test masses and their coatings because of their excellent thermal and optical properties at low temperatures. However, birefringence owing to local impurities and inhomogeneities in the crystal can degrade the performance of the detector. The birefringence measurement or mapping over a two-dimensional area is thus important. This study describes a method for fast birefringence measurements of a large sample by simply combining a series of transmission wavefront error measurements using linearly polarised light with Fizeau interferometers. Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses with a diameter of 22 cm was fully reconstructed. The birefringence information was then used to calculate the transverse beam shapes of the light fields in orthogonal polarisation directions when passing through the substrate. It was possible to find a calculated beam shape consistent with in-situ measurements using the KAGRA interferometer. This technique is crucial for birefringence characterisation of test masses in future detectors, where even larger sizes are used.
@@ I Introduction
Thermal noise is one of the main noise sources for gravitational wave detectors [1, 2]. To reduce thermal noise, future detectors such as the Einstein Telescope [3, 4] and LIGO Voyager [5] concept are designed to operate at cryogenic temperatures. Cosmic Explorer [6, 7] considers cryogenic operations for upgrades. KAGRA [8, 9, 10, 11] is the only detector operating at cryogenic temperatures and pioneers this technique among the current large-scale ground-based detectors. All these detectors use large, high-quality mirrors as their test masses, forming optical resonators to sense the tiny length fluctuations introduced by gravitational waves. Fused silica is commonly used to make test masses at room temperature; however, it is not compatible with cryogenic detectors due to its large mechanical loss at low temperatures. Crystalline materials are considered to be the most promising materials for fabricating cryogenic test masses. Sapphire was selected as the material for KAGRA’s test masses [12] because of its excellent thermal and optical properties at low temperatures [13, 14]. However, recent studies have shown that the birefringence of crystalline substrates and coatings can degrade detector performance [15, 16, 17, 18, 19, 20]. The birefringence fluctuates spatially in magnitude and direction throughout the mirror substrate. Therefore, it is necessary to measure the position dependence of the birefringence over a large area to assess its impact on optical performance.
@@
There are several studies on two-dimensional birefringence measurements of transparent materials [15, 16, 21, 22, 23, 24, 25, 26]. These measurements use either a single-beam scanning technique or expand the beam to a large aperture for imaging using a sensor array. The scanning method is rather time-consuming for large sample sizes, for example, mirrors with diameters usually greater than 20 cm for ground-based gravitational wave detectors, if one wants to obtain good spatial resolution [27, 28]. A phase-shifting method in digital interferometry has been proposed for fast 2D birefringence measurements with high accuracy [29, 30, 31]. Fizeau interferometry is among the most accurate methods for extracting the phase information of an optical component (reflection or transmission phase) over a large area. The Transmission Wavefront Error (TWE) is often used to measure the refractive index homogeneity of a substrate under test. However, in the presence of birefringence a simple TWE measurement does not reveal all the information about the substrate.
@@
In this study, we describe a method for combining a series of TWE map measurements using linearly polarised light in an appropriate way to fully reconstruct refractive index information including birefringence, which is introduced in Section II. Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses (ITM) is extracted, described in Section III. This birefringence information was then used to calculate the transverse beam shape of the light fields in orthogonal polarisation directions when passing through the substrate, which is compared with the in-situ beam shape measurements. Conclusions and discussions are summarised in Section IV.
@@ II Generating a birefringence map with a Fizeau interferometer
Figure 1 shows a schematic diagram of the Fizeau interferometer. We analyse the interference pattern between the light reflected by the reference plane 1 ($E_{\mathrm{1}}$) and the light reflected by the reference plane 2 ($E_{\mathrm{2}}$) on the CCD. By comparing the interference between reference planes 1 and 2, we obtained information on the optical thickness of the sample. Because the optical thickness is the physical thickness multiplied by the index of refraction, by measuring the surface figure maps of S1 (surface with anti-reflection coating) and S2 (surface with highly reflective coating) precisely and subtracting the variation in the physical thickness of the sample, we can extract the fluctuation of the index of refraction in the substrate.
@@ II.1 Transmission of a linearly polarised light through a birefringent medium
where $M$ denotes the round-trip Jones matrix of the substrate. The indices of refraction for the ordinary and extraordinary axes are denoted as $n_{\mathrm{o}}$ and $n^{\prime}_{\mathrm{e}}$ [27], respectively. Then, we can compute the round- trip phase changes of the optical fields polarised along the axes:
@@
$\alpha_{\mathrm{e}}\equiv 2\pi\frac{2dn^{\prime}_{\mathrm{e}}}{\lambda},$ (4)
@@
Using $\alpha_{\mathrm{o}}$ and $\alpha_{\mathrm{e}}$, we define the common and differential phase changes as follows:
$\alpha_{+}\equiv\frac{\alpha_{\mathrm{o}}+\alpha_{\mathrm{e}}}{2}=\frac{4\pi d}{\lambda}(n_{\mathrm{o}}+n^{\prime}_{\mathrm{e}}),$ (5)
$\alpha_{-}\equiv\frac{\alpha_{\mathrm{o}}-\alpha_{\mathrm{e}}}{2}=\frac{4\pi d}{\lambda}(n_{\mathrm{o}}-n^{\prime}_{\mathrm{e}}).$ (6)
@@ II.3 Extraction of $\alpha_{-}$ and $\theta$ from several TWE maps
We can see $\psi(\eta)$ is a function of $\eta$ with parameters $\theta$ and $\alpha_{-}$ determining its shape. Therefore, if we have several measurements of $\psi(\eta)$ with different values of $\eta$, we can use a non-linear fitting algorithm to fit the measured values with Equation (12) to obtain the best fit values of $\theta$ and $\alpha_{-}$.
@@
III Birefringence characterisation of KAGRA sapphire Input test masses
@@
Sapphire is a uniaxial crystal with an optical axis called the c-axis that has no intrinsic birefringence. KAGRA’s sapphire test masses are fabricated such that the input surface is perpendicular to the c-axis. Light propagating in the test mass substrate travels along the c-axis, and theoretically, should experience no birefringence. However, local impurities and inhomogeneities in the crystal, as well as mechanical and thermal stress (e.g. due to the suspension system supporting the mirror), can cause birefringence. These local defects distorted the polarisation distribution of the light field in the transverse plane.
Gravitational wave detectors have very strict requirements on the TWE of light going through input test mass substrates. To achieve this requirement and minimise the influence of crystal inhomogeneity, an ion beam figuring technique was used during the polishing process, in which the thickness of the substrate was precisely adjusted over the surface to ensure that the TWE rms was within a few nanometres [19]. However, this technique optimizes the TWE for the beam with a certain polarisation state, and it cannot reduce birefringence.
@@ III.1 TWE of KAGRA sapphire test mass
The sapphire ITMs have already been installed in KAGRA, and direct birefringence mapping is impossible now. However, characterisation of the sapphire test masses was studied in [33]. The TWE maps of the ITMs were measured several times with linearly polarised light by rotating the sample at different angles, which was equivalent to the measurements obtained by rotating the input polarisation. The TWE maps were measured at the angles of 0, 45, 90 and 135. The methods described in Section II require a rigid angular control. However, the main motivation of the measurements in [33] was to see the stress induced change in $\alpha_{+}$, not $\alpha_{-}$, by different mounting orientations. Therefore, the rotation angle was not precisely controlled. To mitigate the problem, we compared the locations of the markers on the edge of the mirrors in different maps. In this way, we were able to recover the real rotation angle (Figure 3). The TWEs of the KAGRA current ITMs measured with linearly polarised light are around 5 to 6 times larger than the specifications because the vendor used circularly polarised light to measure the path length difference and calculate the removal from surface S2 during the ion beam figuring process (rms values are listed in Table 1). As shown in Figure 3, the 0 map and the 88 map have almost opposite values (the same is true for 49 and 130 maps). By summing the two maps, we can derive the TWE for circularly polarised light, whose rms magnitude is below 5 nm and matches with the specification.
@@ III.2 Constructing birefringence maps from TWE
The biggest challenge for the birefringence characterisation of KAGRA’s current ITMs is that the previous TWE measurement setup was not optimised for birefringence studies. According to Equation (20) in Section II.3, to extract the birefringence information, we need to subtract two pairs of TWE maps and then divide one by the other. In typical Fizeau interferometer measurements, low-order effects such as the piston and surface tilt are removed from the results. When doing measurements with different $\theta$, the distance and alignment between the sample and reference mirror have changed after rotating the sample. This implies that the difference between the two maps contains an unknown piston and tilt. The numerator and denominator in Equation (20), which contain these unknown terms, provide incorrect birefringence information.
The centre of the map also changed after the sample was rotated. However, we observed that the decentring was very small (less than 2 mm), and its effect was negligible compared with the effects of the piston and tilt. In addition, astigmatism was observed on the original TWE maps. The left plot in Figure 4 shows the TWE map after removing the curvature. An astigmatism term can be observed on the map plotted on the right. We observed that the rotation of this astigmatism did not change when the samples were rotated. Therefore, we believe that the astigmatism originates from the gravity-induced deformation of the sample and the astigmatism of the reference mirror used for the Fizeau interferometer. Astigmatism was individually removed from each TWE map.
@@ III.3 Comparison with in-situ measurements
The derived $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY are shown in Figure 10. From these maps, we can see that the birefringence in KAGRA’s current ITMs is inhomogeneous. $\theta$ fluctuates a lot in the range of ±$\pm 50^{\circ}$ due to the fluctuation of the c-axis. The one-way differential phase (retardation) $\alpha_{-}/2$ is in the order of 100 nm, which corresponds to $\Delta n=n_{\mathrm{o}}-n^{\prime}_{\mathrm{e}}=6.7\times 10^{-7}$ (with the thickness of the mirror $d=0.15$ m). If we compare these maps with the beam shape in p-polarisation shown in Figure 8, it can be found that the beam shape is mainly determined by the shape of $\theta$ map. This is because $\alpha_{-}$ is small in our case. The $\sin 2\theta$ term plays a more important part than $\sin\alpha_{-}$ in the amplitude of the field in p-polarisation from Equation (7).
@@ Acknowledgements.
We would like to thank Shinji Miyoki and Tomotada Akutsu at KAGRA for their support for the measurements. We also thank Masaki Ando, Kentaro Somiya, Kentaro Komori and Satoru Takano for insightful discussions. This work was supported by JSPS KAKENHI Grant No. JP20H05854, No. 22H01228, No. 23H01205, No. 23K19053, No. 24K00649, and by JST PRESTO Grant No. JPMJPR200B. KAGRA is supported by MEXT, JSPS Leading-edge Research Infrastructure Program, JSPS Grant-in-Aid for Specially Promoted Research 26000005, JSPS Grant-inAid for Scientific Research on Innovative Areas 2905: JP17H06358, JP17H06361 and JP17H06364, JSPS Core-to-Core Program A. Advanced Research Networks, JSPS Grantin-Aid for Scientific Research (S) 17H06133 and 20H05639 , JSPS Grant-in-Aid for Transformative Research Areas (A) 20A203: JP20H05854, the joint research program of the Institute for Cosmic Ray Research, University of Tokyo, National Research Foundation (NRF), Computing Infrastructure Project of Global Science experimental Data hub Center (GSDC) at KISTI, Korea Astronomy and Space Science Institute (KASI), and Ministry of Science and ICT (MSIT) in Korea, Academia Sinica (AS), AS Grid Center (ASGC) and the National Science and Technology Council (NSTC) in Taiwan under grants including the Rising Star Program and Science Vanguard Research Program, Advanced Technology Center (ATC) of NAOJ, and Mechanical Engineering Center of KEK.